Youve observed the following returns on Yamauchi Corporations stock over the past five years: -29.7 percent, 16.8 percent,
Answer #1

Arithmetic average return=Total return/Total time period

=(-29.7+16.8+36.6+3.9+22.9)/5=10.1%

Return (Return-Average Return)^2
-29.7 (-29.7-10.1)^2=1584.04
16.8 (16.8-10.1)^2=44.89
36.6 (36.6-10.1)^2=702.25
3.9 (3.9-10.1)^2=38.44
22.9 (22.9-10.1)^2=163.84
Total=2533.46%

Standard deviation=[Total (Return-Average Return)^2/(Time Period-1)]^(1/2)

=[2533.46/(5-1)]^(1/2)

=[2533.46/4]^(1/2)

=25.17%(Approx).

Variance=Standard deviation^2

=0.063337(Approx).